OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units.

OpenStudy (anonymous):

put \[x^2+4x-5=7\] and solve for x via \[x^2+4x-12=0\] \[(x+6)(x-2)=0\] \[x+6=0\] \[x=-6\] \[x-2=0\] \[x=2\] and in this problem the negative answer makes no sense so it is just x = 2

jhonyy9 (jhonyy9):

x*2 +4x -5 = 7*2 --- x*2 +4x -5 =49 --- x*2 +4x -54 =0 4 +/- (16 + 216)*(1/2) 4 +/- (232)*(1/2) 2(2+/- (58)*(1/2) x_1,_2 = --------------------- = ----------------- = ---------------- = 2 2 2 =2+/- (58)*(1/2) x_1,_2 = 2+/- radical 58